OpenStudy (anonymous):

i think a Which answer shows this equation written in standard form? 9 – 4(x – y) = –6x + 6 A. 2x + 4y = 3 B. 2x + 4y = –3 C. 2x – 4y = 3 D. –10x + 4y = –3

OpenStudy (susanna):

9 – 4(x – y) = –6x + 6 9-4x+4y=-6x+6 9-10x+4y=6 -10x+4y=-3